calculate the number of molecules and number of atoms present in 5.60l of ozone at n.t.p
1
Expert's answer
2012-05-08T08:57:46-0400
n = V/22.4 = 0,25 mol

in 1 mol is 6,02x10^23 (Avogadro constant)
in 0,25 mol is 1.505x10^23
number of molecules :1.505x10^23
In ozone(O3) are 3 atoms of oxygen,so number of atoms :3*1.505x10^23 =4,515x10^23